OpenStudy (anonymous):

A gas has a volume of 590 mL at a temperature of –55.0 C. What volume will the gas occupy at 30.0 C?

OpenStudy (aaronq):

use: \(\dfrac{V_1}{T_1}=\dfrac{V_2}{T_2}\) (this is charles's law) ps. the temperature has to be in Kelvin

OpenStudy (anonymous):

Someone told me Since V1T1 = V2T2 (590mL)(-550C) divided by (30 0C) = 1081,67 mL but I don't know how to get or convert the answer to Kelvin units...

OpenStudy (aaronq):

to convert to K from celsius just add 273.15 .. so -55 celsius to kelvin 273 + (-55)= 218 K

OpenStudy (anonymous):

Oooh.. Is that the final product?

OpenStudy (aaronq):

that would be one of the temperatures given in Kelvin, but it's not the answer. you'd still have to convert the the other temp and solve for the new volume

OpenStudy (anonymous):

Oh gosh... how would I do that?

OpenStudy (aaronq):

to convert from celsius to kelvin simply add 273 to the temperature. e.g. \(5^oC \) to kelvin.. 5 + 273 = 278 K use the equation i wrote above (in my first post) to solve for the new volume

OpenStudy (anonymous):

590/218 = V2/303 V2 = 590 (303/218) = 820 ml

OpenStudy (aaronq):

thats right \(\checkmark\)

OpenStudy (jtfan2000):

(590mL)(-55C) divided by (30C) = 1081.67mL